OpenStudy (anonymous):

Simplify: 21 - 3 × 6 + 32 the choices are ...... a.21 b.12 c. 48 d. 117

OpenStudy (anonymous):

35

OpenStudy (anonymous):

21 - 18 + 32 = 3 + 32 = 35

OpenStudy (anonymous):

erm um the chocies are 21 12 48 117

OpenStudy (anonymous):

PEMDAS google it

OpenStudy (anonymous):

yea I know please my dear aunt sally I don't get it!

OpenStudy (anonymous):

35 is not one of the chocies

OpenStudy (anonymous):

Then you either copied the question incorrectly or the choices are wrong. 35 is definitely correct

OpenStudy (anonymous):

Yeah lol. The answer IS 25.

OpenStudy (anonymous):

35

OpenStudy (anonymous):

erm the question is21 - 3 × 6 + 32 and 2 above 3

OpenStudy (anonymous):



OpenStudy (anonymous):

not